OpenStudy (anonymous):

Factor completely: 2x^3 + 4x^2 + 6x + 12

OpenStudy (anonymous):

Factor completely: x^3 – 2x2 + 6x – 12 = x^2(x-2) + 6(x-2) =(x-2)(x^2+6) ------------------ . Factor completely: x^2 + 17x + 16 = (x+16)(x+1) ------------------- Factor completely: 8x^2 + 6x – 5 = 8x^2+10x-4x-5 = 2x(4x+5)-(4x+5) = 4x+5)(2x-1) x^3 – 2x^2 + 6x – 12 x^2(x-2)+6(x-2) Factoring by grouping. (x^2+6)(x-2) . x^2 + 17x + 16 (x+16)(x+1) . 8x^2 + 6x – 5 8*-5=-40 What 2 factors -40 when added equal +6? Ans: 10, -4 8x^2+10x-4x-5 Put -4x and 10xin the middle in either order. 2x(4x+5)-(4x-5) Factoring by grouping again. (2x-1)(4x+5) .